I am stuck on correct answer for:
if the sum of 3 consecutive integers is 36, what is the smallest of these integers?
I believe answer is 11 but I need to verify my work. Can you show me the breakdown to get correct resolve?

Answer by stanbon(48569)

if the sum of 3 consecutive integers is 36, what is the smallest of these integers?
--------------
Let the consecutive integers be x, x+1, x+2
EQUATION:
x+x+1+x+2=36
3x+3=36
x+1=12
x=11
x+1=12
x+3=13
Cheers,
Stan H.